OpenStudy (anonymous):

Help Please! <3 Will Give Medal!(: 1. Anna burned 15 calories per minute running x minutes and 10 calories per minute hiking for y minutes. She spent a total of 60 minutes running and hiking and burned 700 calories. The system of equations shown below can be used to determine how much time Anna spent on each exercise. 15x + 10y = 700 x + y = 60 what is the value of x, the minutes Anna spent running? a. 10 b. 20 c. 30 d. 40

OpenStudy (anonymous):

http://www.purplemath.com/modules/systlin5.htm

OpenStudy (nikato):

use substitution subtract x on both sides in the second equation x+y=60 -x -x ----------- y=60-x then substitute this into the first equation 15x+10(60-x)=700 now distribute and then solve for x